# Microphone Array Geometry Independent Multi-Talker Distant ASR: NTT System for the DASR Task of the CHiME-8 Challenge

### v1

@@ Abstract
In this paper, we introduce a multi-talker distant automatic speech recognition (DASR) system we designed for the DASR task 1 of the CHiME-8 challenge. Our system performs speaker counting, diarization, and ASR. It handles various recording conditions, from diner parties to professional meetings and from two to eight speakers. We perform diarization first, followed by speech enhancement, and then ASR as the challenge baseline. However, we introduced several key refinements. First, we derived a powerful speaker diarization relying on end-to-end speaker diarization with vector clustering (EEND-VC), multi-channel speaker counting using enhanced embeddings from EEND-VC, and target-speaker voice activity detection (TS-VAD). For speech enhancement, we introduced a novel microphone selection rule to better select the most relevant microphones among the distributed microphones and investigated improvements to beamforming. Finally, for ASR, we developed several models exploiting Whisper and WavLM speech foundation models. We present the results we submitted to the challenge and updated results we obtained afterward. Our strongest system achieves a 63% relative macro tcpWER improvement over the baseline and outperforms the challenge best results on the NOTSOFAR-1 meeting evaluation data among geometry-independent systems.
@@ 1 Introduction
Within the last decade, rapid progress in deep learning has led to dramatic improvement in automatic speech recognition (ASR) performance. However, it remains challenging to recognize speech in multi-talker recordings, especially when using distant microphones [1]. The problem of multi-talker distant ASR (DASR) consists not only of transcribing what was spoken (i.e., ASR) but also correctly estimating who uttered the utterances and when (i.e., speaker diarization) [2, 3]. Besides, it should be robust to overlapping speech, background noise, and reverberation, which are common distortions observed when recording with distant microphones in everyday living and working environments. This makes the multi-talker ASR problem extremely challenging, but solving it would have great implications, as it would allow broadening the range of ASR applications to automatic meeting minute generation, smart human-robot interactions, etc.
@@
The CHiME challenge series has been a leading activity in triggering research in multi-talker distant ASR (DASR). It originated with the CHiME-1-2 challenges [4, 5], which were concerned with DASR in living rooms. At that time, the problem consisted of recordings of simple commands or read sentences from a single speaker recorded by a distant microphone array in the presence of various noise sources that can appear in a living room, such as a vacuum cleaner, children’s voices, etc. CHiME-3-4 [6] extended the problem to cover four different recording conditions, including bus, café, street, and pedestrian. CHiME-5-6 [7, 8] increased the complexity by switching to multi-talker conversations in a dinner party setting recorded with distributed microphone arrays. This also added the requirement to perform speaker diarization. More recently, the DASR tasks of CHiME-7-8 challenges (task 1) [2, 3] further increased the complexity by requiring building a single array geometry independent system that could perform multi-talker DASR for 1) diverse recording scenarios (dinner parties, interviews, meetings), 2) recordings with various distributed microphone array configurations, 3) with a number of speakers varying from two to eight, 4) recording durations from few minutes to more than one hour, and 5) limited amount of in-domain training data. CHiME-8 challenge also had another DASR task, called NOTSOFAR-1 task (task 2) [9], which focused on DASR of meetings of four to eight speakers recorded with a single microphone array, with known array geometry, in relatively well-controlled conditions consisting of office meeting rooms. DASR and NOTSOFAR-1 ran as separate tasks, but the NOTSOFAR-1 data was also used as one of the datasets in the DASR task 1 evaluation. In this paper, we describe the system we proposed for the DASR task of the CHiME-8 challenge [3].
@@
There are different ways to implement a multi-talker DASR system to perform both ASR and diarization while being robust to noise, reverberation, and overlapping speech. Here, we briefly review the two main approaches [10], which vary by the processing order of the diarization, speech separation/enhancement, and ASR modules, i.e., separation-first and diarization-first approaches.
@@
The continuous speech separation (CSS) pipeline [11, 12, 10] is an example of a separation-first approach, which performs speech separation on short time segments, then transcribes each segment, and finally performs diarization by clustering speaker embeddings computed on the separation outputs. The advantage of this approach is that ASR and diarization are greatly simplified if the speech signals can be well separated beforehand. For example, after separation, we can obtain speaker embedding less affected by overlapping speech, which can reduce speaker confusion of diarization. Besides, diarization can exploit the ASR output to, e.g., refine speaker activity prediction [13]. Such a separation-first pipeline relies heavily on having good separation pre-processing, typically implemented with a combination of neural-network (NN) -based separation and mask-based beamformer. The separation-first pipeline was used as a baseline for the NOTSOFAR-1 task 2 of the CHiME challenge. In that meeting recognition scenario, the recording conditions are relatively well-controlled, making NN-based separation very effective. It is, however, challenging to use such a separation-first pipeline for the domain- and microphone array geometry-independent DASR systems, such as required for the DASR task 1 of CHiME-8, where the recording conditions are much more diverse, making it difficult to train a robust NN-based separation frontend.
@@
The NeMo baseline of the DASR task 1 of the CHiME-8 challenge is an example of a diarization-first system [3]. The processing flow begins with diarization, followed by guided source separation (GSS) [14] for speech enhancement (SE) and then ASR. The idea behind this approach is that it is easier to build a robust diarization system than separation because diarization requires estimating coarser information, i.e., the speech activity (discrete values) for diarization compared to waveforms for separation. Then, robust speech separation can be performed using GSS, which consists of an adaptive speech separation method such as complex angular central Gaussian mixture model (cACGMM) [15] conditioned on the speaker activity prediction of the diarization pre-processing. Such a pipeline has been employed by most systems proposed for the CHiME-6, 7, and 8 [16, 17, 18, 19, 20, 21, 22] because it can be used with distributed microphone arrays and is relatively robust to the very challenging recording conditions for which NN-based separation remains challenging.
@@
In this paper, we propose a modification of the diarization-first pipeline, which can partially utilize the advantage of the separation-first pipeline. Current state-of-the-art diarization systems, such as end-to-end diarization with vector clustering (EEND-VC) [23, 24, 25], usually rely on a two-step processing, i.e., local segmentation and global clustering. The first step, local segmentation, consists of predicting the speech activity (i.e., when) of each speaker in short audio segments of up to, e.g., 15-80 sec, assuming that the number of locally active speakers remains below a pre-set maximum number of local speakers, e.g., four. We then perform clustering of the speaker embeddings associated with each speaker activity in each segment to associate the local speaker activity with a recording-level speaker label (i.e., who). This allows us to assemble the activity of each speaker together to form recording-level diarization results despite possible speaker ordering ambiguity between segments. It also enables handling a global number of speakers superior to the maximum number of local speakers. In this paper, we exploit another advantage of the EEND-VC approach. Because the segmentation and clustering can be performed independently, we propose to perform GSS-based separation between these two steps, as GSS only requires local segmentation information. This approach, which we call segmentation-first, can bring some of the benefits of the separation-first pipeline, i.e., extracting speaker embeddings on enhanced speech improves their quality, resulting in better clustering accuracy.
Figure 1: Overview of the proposed multi-talker DASR system.
@@
Figure 1 shows an overview of our proposed multi-talker DASR system. It consists of the modified EEND-VC frontend, which inserts GSS between segmentation and clustering. For clustering, we set the number of clusters to the number of speakers estimated using a novel multi-channel speaker counting approach. We then refine the diarization results using target speaker voice activity detection (TS-VAD) [26] implemented based on the memory-aware multi-speaker embedding with sequence-to-sequence architecture (NSD-MS2S) [27], which achieved top performance in the CHiME-7 challenge [17]. For SE, we enhance the target speaker for each segment using GSS combined with weighted prediction error (WPE) [28, 29] and spatial-prediction multichannel Wiener filter (SP-MWF) beamformer [30, 31]. In addition, we designed a novel microphone subset selection approach to reduce the influence of unreliable microphone signals. Finally, we perform ASR using four different backends based on Whisper, NeMo, and a WavLM-based RNN-T model. To obtain optimal results, we performed language model re-scoring and combined the output of the four systems using recognizer output voting error reduction (ROVER) -based system combination [32].
@@
    Handling various numbers of speakers and recording durations: Our diarization system relies on EEND-VC, which can handle an arbitrary number of speakers. With EEND-VC, the result of the clustering of the speaker embeddings provides the speaker counts. However, the counting accuracy can worsen when speaker embeddings are unreliable or when there are not enough embeddings when the recordings are short. To improve speaker counting accuracy, we thus propose to decouple the clustering for speaker counting from the clustering for assigning speaker labels. We increase the reliability of the estimated speaker counts by increasing the number of embeddings for clustering, computing them on shorter segments, and aggregating the results over the different microphones. Then, we can perform speaker embedding clustering to attribute speaker labels by fixing the number of clusters to the estimated speaker counts.
@@
    Dealing with insufficient and unreliable transcriptions for training: One issue with the CHiME-8 DASR dataset is that the amount of reliable training data for ASR is limited. Using pre-trained foundation models can partially alleviate this issue, but the best performance is obtained when these models are finetuned on the target task. We thus propose two approaches to deal with data scarcity. First, we propose augmenting the training data with un-transcribed data from VoxVeleb 1 and 2 datasets. We apply contrastive data selection [35] to select data similar to the CHiME-8 task 1 datasets’ conditions. In addition, when retraining the Whisper model, we introduce a curriculum learning scheme, which filters out training samples with many recognition errors.
This paper extends our challenge system description paper [21] and our conference papers describing the diarization part [36, 37]. Compared to our previous publications, we provide more exhaustive descriptions of all parts of our system, especially the diarization and SE frontends, and include extensive experimental analysis and ablation studies to better understand each module’s contribution. Moreover, we provide new post-challenge results that further improved our performance, narrowing the gap with the top CHiME-8 system [20] and surpassing it on the NOTSOFAR-1 data.
In the remainder of the paper, we provide an overall description of our proposed multi-talker DASR system in Section 2. We describe diarization and speaker counting in Section 3, SE in Section 4, and the ASR backend in Section 5. We then provide experimental results and analysis in Section 7, with extensive ablation for the different modules in A. Finally, we conclude the paper in Section 8.
This paper attempts to provide an exhaustive description of our system, which makes the overall paper rather long. Readers interested in the details of each module can read the related sections 3, 4 and 5. Readers, mostly interested in the experiment analysis, can read the overall description in Section 2 and then proceed directly to the experiments of Section 7.
@@ 2 Overview of proposed system
The multi-talker DASR consists of transcribing a multi-talker conversation captured with distant microphones. Here, transcribing also includes assigning a speaker label and the boundaries for each utterance (start and end time). [^1] Note that the system does not provide absolute speaker identities but should only associate a consistent speaker label for each speaker in a recording. We can model the signal received at a distant microphone as:
@@
where $\mathbf{y}_{m}$, $\mathbf{x}_{m,s}$ and $\mathbf{n}_{m}\in\mathbb{R}^{T}$ are the noisy observation, speech, and noise signals, respectively. $m$ and $s$ are the microphone and speaker indexes, respectively. $S$ represents the number of speakers and $T$ the signal duration. For simplicity of the notation, we did not consider reverberation in the notations. We define the set of all microphone signals as $\mathbf{Y}=\{\mathbf{y}_{m}\}_{1\leq m\leq M}$, where $M$ is the number of microphones. In the CHiME-8 DASR task, the number of speakers, $S$, is considered unknown and varies depending on the recordings from two to eight. The number of microphones, $M$, also varies depending on the datasets, and the array geometry is unknown.
The speech signal $\mathbf{x}_{m,s}$ consists of a long signal containing all utterances spoken by speaker $s$ in the recording and zeros when the speaker is not active. The DASR system aims to output transcriptions and timings of all utterances of all speakers in the recording. Let us define the output $\mathbf{O}=\{\mathbf{O}_{s}\}_{1\leq s\leq S}$ as the sets of all utterances of all speakers, where $\mathbf{O}_{s}=\{\mathbf{o}_{s,j}\}_{1\leq j\leq J_{s}}$ represents the set of all $J_{s}$ utterances associated with speaker $s$. Here, $\mathbf{o}_{s,j}=(\mathbf{v}_{s,j},\mathbf{b}_{s,j})$ is the $j$-th utterance of speaker $s$, where $\mathbf{v}_{s,j}$ is the transcription or token sequence, $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$ represents the utterance boundaries, and $t^{\text{start}}_{s,j}$ and $t^{\text{end}}_{s,j}$ are the start and end times of the utterance, respectively.
The DASR system estimates $\mathbf{O}$ given the microphone observations as,
@@
We implement the proposed multi-talker DASR system as a cascade of diarization, SE, and ASR modules, as shown in Fig. 1. Below, we briefly define each module to emphasize how they are interconnected. In the subsequent sections, we explain the details of each module.
Diarization: The diarization module operates on the multi-channel signals of the whole recordings. It predicts the speaker activity for each speaker in the recording. Here, by abuse of notations, we consider an estimated active speaker segment as an utterance, although it may not match an actual reference utterance. The diarization module estimates thus the utterance boundaries and assigns a speaker label to each estimated utterance as,
@@
Speech enhancement: The SE module aims at isolating the speech signal of the target speaker in each utterance, removing interference, noise, and reverberation. It operates on each utterance found by the diarization module and is conditioned on the estimated speakers’ activity information. We can formulate its processing as,
@@
ASR: Finally, we perform ASR on each enhanced signal channel speech signal, $\hat{\mathbf{x}}_{s,j}$, to obtain the transcriptions $\hat{\mathbf{v}}_{s,j}$ as,
@@
where $\text{ASR}(\cdot)$ represents the ASR module. To improve performance, we perform ASR using several models and perform system-combination using ROVER. In addition, we also perform language model rescoring. We provide the details of the ASR module in Section 5.

### v2

# Microphone Array geometry-independent Multi-Talker Distant ASR: NTT System for DASR Task of the CHiME-8 Challenge

@@ Abstract
In this paper, we introduce a multi-talker distant automatic speech recognition (DASR) system we designed for the DASR task 1 of the CHiME-8 challenge. Our system performs speaker counting, diarization, and ASR. It handles a variety of recording conditions, from dinner parties to professional meetings and from two speakers to eight. We perform diarization first, followed by speech enhancement, and then ASR as the challenge baseline. However, we introduced several key refinements. First, we derived a powerful speaker diarization relying on end-to-end speaker diarization with vector clustering (EEND-VC), multi-channel speaker counting using enhanced embeddings from EEND-VC, and target-speaker voice activity detection (TS-VAD). For speech enhancement, we introduced a novel microphone selection rule to better select the most relevant microphones among those distributed microphones and investigated improvements to beamforming. Finally, for ASR, we developed several models exploiting Whisper and WavLM speech foundation models. In this paper, we present the original results we submitted to the challenge and updated results we obtained afterward. Our strongest system achieves a 63% relative macro tcpWER improvement over the baseline and outperforms the challenge best results on the NOTSOFAR-1 meeting evaluation data among geometry-independent systems.
@@ 1 Introduction
Within the last decade, rapid progress in deep learning has led to dramatic improvements in the performance of automatic speech recognition (ASR). However, it remains challenging to recognize speech in multi-talker recordings, especially when using distant microphones [1]. The problem of multi-talker distant ASR (DASR) consists of not only transcribing what was spoken (i.e., ASR) but also correctly estimating who uttered the utterances and when (i.e., speaker diarization) [2, 3]. Moreover, it should be robust to overlapping speech, background noise, and reverberation, which are common distortions observed when recording with distant microphones in everyday living and working environments. This makes the multi-talker ASR problem extremely challenging, but solving it would have great implications, as it would enable broadening the range of ASR applications to automatic meeting minute generation, smart human-robot interactions, etc.
@@
The CHiME challenge series has been a leading activity in triggering research on multi-talker distant ASR (DASR). It originated with the CHiME-1-2 challenges [4, 5], which involved DASR in living rooms. At that time, the problem consisted of recordings of simple commands or sentences read from a single speaker recorded by a distant microphone array in the presence of various noise sources that can appear in a living room, such as a vacuum cleaner, children’s voices, etc. The CHiME-3-4 [6] extended the problem to cover four different recording conditions, including bus, café, street, and pedestrian. The CHiME-5-6 [7, 8] increased the complexity by switching to multi-talker conversations in a dinner party setting recorded with distributed microphone arrays. This also added the requirement to perform speaker diarization. More recently, the DASR tasks of the CHiME-7-8 challenges (task 1) [2, 3] further increased the complexity by requiring the construction of a single array geometry-independent system that could perform multi-talker DASR for 1) diverse recording scenarios (dinner parties, interviews, meetings), 2) recordings with various distributed microphone array configurations, 3) recordings with the number of speakers varying from two to eight, 4) recording durations from a few minutes to more than one hour, and 5) limited amounts of in-domain training data. The CHiME-8 challenge also featured another DASR task called the NOTSOFAR-1 task (task 2) [9], which focused on the DASR of meetings of four to eight speakers recorded with a single microphone array, with known array geometry, in relatively well-controlled conditions consisting of office meeting rooms. DASR and NOTSOFAR-1 ran as separate tasks, but the NOTSOFAR-1 data was also used as one of the datasets in the DASR task 1 evaluation. In this paper, we describe the system we proposed for the DASR task of the CHiME-8 challenge [3].
@@
There are different ways to implement a multi-talker DASR system to perform both ASR and diarization while being robust to noise, reverberation, and overlapping speech. Here, we briefly review the two main approaches [10], separation-first and diarization-first approaches, which vary in terms of the processing order of the diarization, speech separation/enhancement, and ASR modules.
@@
The continuous speech separation (CSS) pipeline [11, 12, 10] is an example of a separation-first approach, which performs speech separation on short time segments, then transcribes each segment, and finally performs diarization by clustering speaker embeddings computed on the separation outputs. The advantage of this approach is that ASR and diarization are greatly simplified if the speech signals can be well separated beforehand. For example, after separation, we can obtain speaker embedding less affected by overlapping speech, which can reduce the speaker confusion of diarization. In addition, diarization can exploit the ASR output in order to, e.g., refine speaker activity prediction [13]. Such a separation-first pipeline relies heavily on having good separation pre-processing, typically implemented with the combination of neural-network (NN) -based separation and a mask-based beamformer. The separation-first pipeline was utilized as a baseline for the NOTSOFAR-1 task 2 of the CHiME challenge. In that meeting recognition scenario, the recording conditions are relatively well-controlled, making NN-based separation very effective. It is, however, challenging to use such a separation-first pipeline for the domain- and microphone array geometry-independent DASR systems (which are required for the DASR task 1 of CHiME-8), where the recording conditions are much more diverse, making it difficult to train a robust NN-based separation frontend.
@@
The NeMo baseline of the DASR task 1 of the CHiME-8 challenge is an example of a diarization-first system [3]. The processing flow begins with diarization, followed by guided source separation (GSS) [14] for speech enhancement (SE) and then ASR. The idea behind this approach is that it is easier to build a robust diarization system than separation because diarization requires estimating coarser information, i.e., the speech activity (discrete values) for diarization compared to waveforms for separation. Then, robust speech separation can be performed using GSS, which consists of an adaptive speech separation method such as a complex angular central Gaussian mixture model (cACGMM) [15] conditioned on the speaker activity prediction of the diarization pre-processing. Such a pipeline has been utilized by most systems proposed for the CHiME-6, 7, and 8 [16, 17, 18, 19, 20, 21, 22] because it can be used with distributed microphone arrays and is relatively robust to the very challenging recording conditions for which NN-based separation remains challenging.
@@
In this paper, we propose a modification of the diarization-first pipeline that can partially utilize the advantage of the separation-first pipeline. Current state-of-the-art diarization systems, such as end-to-end diarization with vector clustering (EEND-VC) [23, 24, 25], usually rely on a two-step processing, i.e., local segmentation and global clustering. The first step, local segmentation, consists of predicting the speech activity (i.e., when) of each speaker in short audio segments of up to, e.g., 15–80 sec, assuming that the number of locally active speakers remains below a pre-set maximum number of local speakers, e.g., four. We then perform clustering of the speaker embeddings associated with each speaker activity in each segment to associate the local speaker activity with a recording-level speaker label (i.e., who). This allows us to assemble the activity of each speaker together to form recording-level diarization results despite possible speaker ordering ambiguity between segments. It also enables handling a greater number of global speakers compared to the maximum number of local speakers. In this paper, we exploit another advantage of the EEND-VC approach: name, since the segmentation and clustering can be performed independently, we propose performing GSS-based separation between these two steps, as GSS only requires local segmentation information. This approach, which we call segmentation-first, can harness some of the benefits of the separation-first pipeline, i.e., extracting speaker embeddings on enhanced speech improves their quality, resulting in better clustering accuracy.
Figure 1: Overview of proposed multi-talker DASR system.
@@
Figure 1 shows an overview of our proposed multi-talker DASR system. It consists of the modified EEND-VC frontend, which inserts GSS between segmentation and clustering. For clustering, we set the number of clusters to the number of speakers estimated using a novel multi-channel speaker counting approach. We then refine the diarization results using target speaker voice activity detection (TS-VAD) [26] implemented based on the memory-aware multi-speaker embedding with sequence-to-sequence architecture (NSD-MS2S) [27], which achieved the top performance in the CHiME-7 challenge [17]. For SE, we enhance the target speaker for each segment using GSS combined with weighted prediction error (WPE) [28, 29] and spatial-prediction multichannel Wiener filter (SP-MWF) beamformer [30, 31]. In addition, we designed a novel microphone subset selection approach to reduce the influence of unreliable microphone signals. Finally, we perform ASR using four different backends based on Whisper, NeMo, and a WavLM-based RNN-T model. To obtain optimal results, we performed language model re-scoring and combined the output of the four systems using recognizer output voting error reduction (ROVER) -based system combination [32].
@@
    Handling various numbers of speakers and recording durations: Our diarization system relies on EEND-VC, which can handle an arbitrary number of speakers. With EEND-VC, the result of the clustering of the speaker embeddings provides the speaker counts. However, the counting accuracy can worsen when speaker embeddings are unreliable or if there are not enough embeddings when the recordings are short. To improve speaker counting accuracy, we therefore propose decoupling the clustering for speaker counting from the clustering for assigning speaker labels. We improve the reliability of the estimated speaker counts by increasing the number of embeddings for clustering, computing them on shorter segments, and aggregating the results over the different microphones. Then, we can perform speaker embedding clustering to assign speaker labels by fixing the number of clusters to the estimated speaker counts.
@@
    Dealing with insufficient and unreliable transcriptions for training: One issue with the CHiME-8 DASR dataset is that the amount of reliable training data for ASR is limited. Using pre-trained foundation models can partially alleviate this issue, but the best performance is obtained when these models are fine-tuned on the target task. We propose two approaches to deal with data scarcity. First, we augment the training data with un-transcribed data from the VoxVeleb 1 and 2 datasets and apply contrastive data selection [35] to select data similar to the CHiME-8 task 1 datasets’ conditions. Second, when retraining the Whisper model, we introduce a curriculum learning scheme, which filters out training samples with many recognition errors.
This paper is an extention of our challenge system description paper [21] and our conference papers describing the diarization part [36, 37]. Compared to our previous publications, we provide more exhaustive descriptions of all parts of our system, especially the diarization and SE frontends, and include extensive experimental analysis and ablation studies to clarify each module’s contribution. Moreover, we provide new post-challenge results that further improved our performance, narrowing the gap with the top CHiME-8 system [20] and surpassing it on the NOTSOFAR-1 data.
In the remainder of the paper, we provide an overall description of our proposed multi-talker DASR system in Section 2. We describe diarization and speaker counting in Section 3, SE in Section 4, and the ASR backend in Section 5. We then discuss experimental results and analysis in Section 7, with extensive ablation for the different modules in A. We conclude the paper in Section 8.
As our aim is to provide an exhaustive description of our system, the overall paper is rather long. Readers interested in the details of each module can skip ahead to the relevant parts (diarization in Section 3, SE in Section 4 and ASR in Section 5), and those who are interested in the experiment analysis can read the overall description in Section 2 and then proceed directly to the experiments in Section 7.
@@ 2 Overview of proposed system
The multi-talker DASR consists of transcribing a multi-talker conversation captured with distant microphones. Here, transcribing also includes assigning a speaker label and the boundaries for each utterance (start and end times). [^1] Note that the system does not provide absolute speaker identities but should only associate a consistent speaker label for each speaker in a recording. We can model the signal received at a distant microphone:
@@
where $\mathbf{y}_{m}$, $\mathbf{x}_{m,s}$ and $\mathbf{n}_{m}\in\mathbb{R}^{T}$ are the noisy observation, speech, and noise signals, respectively. $m$ and $s$ are the microphone and speaker indexes, respectively. $S$ represents the number of speakers and $T$ the signal duration. For simplicity, we do not consider reverberation in the notations. We define the set of all microphone signals as $\mathbf{Y}=\{\mathbf{y}_{m}\}_{1\leq m\leq M}$, where $M$ is the number of microphones. In the CHiME-8 DASR task, the number of speakers, $S$, is considered unknown and varies from two to eight depending on the recordings. The number of microphones, $M$, also varies depending on the datasets, and the array geometry is unknown.
The speech signal $\mathbf{x}_{m,s}$ consists of a long signal containing all utterances spoken by speaker $s$ in the recording and is zeros when the speaker is not active. The DASR system aims to output transcriptions and timings of all utterances of all speakers in the recording. Let us define the output $\mathbf{O}=\{\mathbf{O}_{s}\}_{1\leq s\leq S}$ as the sets of all utterances of all speakers, where $\mathbf{O}_{s}=\{\mathbf{o}_{s,j}\}_{1\leq j\leq J_{s}}$ represents the set of all $J_{s}$ utterances associated with speaker $s$. Here, $\mathbf{o}_{s,j}=(\mathbf{v}_{s,j},\mathbf{b}_{s,j})$ is the $j$-th utterance of speaker $s$, where $\mathbf{v}_{s,j}$ is the transcription or token sequence, $\mathbf{b}_{s,j}=[t^{\text{start}}_{s,j},t^{\text{end}}_{s,j}]$ represents the utterance boundaries, and $t^{\text{start}}_{s,j}$ and $t^{\text{end}}_{s,j}$ are the start and end times of the utterance, respectively.
The DASR system estimates $\mathbf{O}$ given the microphone observations, as
@@
We implement the proposed multi-talker DASR system as a cascade of diarization, SE, and ASR modules, as shown in Figure 1. Below, we briefly define each module to demonstrate how they are interconnected. In the subsequent sections, we explain the details of each module.
Diarization: The diarization module operates on the multi-channel signals of all recordings and predicts the speaker activity for each speaker in the recording. Here, by abuse of notations, we consider an estimated active speaker segment as an utterance, although it may not match an actual reference utterance. The diarization module thus estimates the utterance boundaries and assigns a speaker label to each estimated utterance, as
@@
Speech enhancement: The SE module aims to isolate the speech signal of the target speaker in each utterance, removing interference, noise, and reverberation. It operates on each utterance found by the diarization module and is conditioned on the estimated speakers’ activity information. We can formulate its processing, as
@@
ASR: Finally, we perform ASR on each enhanced signal channel speech signal, $\hat{\mathbf{x}}_{s,j}$, to obtain the transcriptions $\hat{\mathbf{v}}_{s,j}$ as
@@
where $\text{ASR}(\cdot)$ represents the ASR module. To improve the performance, we perform ASR using several models and perform system-combination using ROVER. In addition, we also perform language model rescoring. We provide the details of the ASR module are provided in Section 5.